3. The uncoiling of DNA is a first order reaction. Its activation energy is 420 kJat 37oC, the rate constant is 4.90 X 10 mind What is the half-life of the uncoiling if the organism has a temperature of 370C (104F). Answer #3
Answer #1

Given the reaction is first order.

So half life,t1/2=0.693/k

Where

K=rate constant=4.90*10^-4 min^-1

So t1/2=0.693/k =1414 min